OpenStudy (anonymous):

There are two consecutive even numbers such that five times the smaller minus three times the greater is 6. What is the smaller number?

OpenStudy (anonymous):

\[5(x-2)-3x=6\] is your equation.

OpenStudy (anonymous):

solve via \[5x-10-3x=6\] \[2x-10=6\] \[2x=16\] \[x=8\]

jhonyy9 (jhonyy9):

6 8

OpenStudy (anonymous):

bigger number is 8, smaller is 8-2=6

jhonyy9 (jhonyy9):

those are